### v1

# Neuronal Spike Trains as Functional-Analytic Distributions: Representation, Analysis, and Significance

Abstract
The action potential constitutes the digital component of the signaling dynamics of neurons. But the biophysical nature of the full time course of the action potential associated with changes in membrane potential is fundamentally and mathematically distinct from its representation as a discrete set of events that encode when action potentials triggered in a collection spike trains. In this paper, we rigorously explore from first principles the transition and modeling from the standard biophysical picture of a single action potential to its representation as a spike in a spike train. In particular, we adopt a functional-analytic framework, using Schwartz distribution theory to represent spike trains as generalized Dirac delta functions acting on smooth test functions. We then show how and why this representation transcends a purely descriptive formalism to support deep downstream analysis and modeling of spike train neural dynamics in a mathematically consistent way.
1 Introduction
An action potential, as a dynamic transient change in the membrane potential of a neuron, has the shape and kinetics it has because of the underlying biophysics of the change in membrane permeability (conductance) to (primarily) $Na^{+}$ and $K^{+}$ ion channels. The conditions responsible for the diffusion (concentration) and electrochemical driving forces that move ions across the membrane are necessarily set up before an action potential is triggered, resulting in the instantaneous passive flux of $Na^{+}$ and $K^{+}$ ions acting as charge carriers responsible for the currents and resultant transient changes in membrane potential we call an action potential.
We assume the reader is familiar with (at least the qualitative) thermodynamic and biophysical description of the action potential summarized in the preceding paragraph. To varying degrees of detail, any standard textbook in neurophysiology can be consulted. Vander’s Human Physiology is particularly good at an introductory undergraduate level [widmaier2023vanders], while Kendel et al. is the standard graduate-level textbook in the field [kandel2021principles]. Although there are many other superb resources, both online and in print.
Under normal physiological conditions, the shape of the standard neuronal action potential is stereotyped and unchanging. Every time an action potential is triggered, it proceeds to completion uninterupted in exactly the same way with the same time course, independent of anything else the neuron is doing. As alluded to above, this is because the biophysical conditions responsible for the processes that mitigate the changes in membrane potential are set a priori. From an information-theoretic perspective, the action potential constitutes the digital component of neurons’ signaling dynamics. The information it contributes to the system is in the timing of its occurrence, not in its shape, amplitude, or time course.
It is tempting to think of a spike as a very sharp voltage waveform. A tall, narrow pulse that looks like a spike on a compressed time axis, compared to the time axis of a full action potential (Figure 1). But a spike is not a compressed action potential, nor does it encode or represent a change in the membrane potential. In fact, a spike is not a signal in the usual sense at all.
Figure 1: The often shown but inacurate picture of an action potential, which is a stereotyped change in the membrane potential due to underlying biophysics, compressed to a spike intended to represent it.
At the level of spiking models, what is recorded is not waveforms but events, the times at which something happened. The spike encodes the information communicated by the action potential being triggered, not the action potential itself. For example, a neuron fires at 12.3 ms. Then again at 18.7 ms. Then at 41.2 ms. The spike train of that neuron, at its core, is just a collection of time stamps.
This immediately creates a disconnect with the mathematical machinery we typically use to describe and model physical systems. Differential equations, filters, convolutions, and dynamical systems are built to operate on functions of time: objects that assign a value to every instant. A list of isolated time points does not fit naturally into this framework. Critically, we need a way to mathematically model spike trains that maintains functional consistency. We need a description of spike trains that represents them for what they are, but in a way that supports interaction with functions and operations that model physical processes that are very different mathematically from the description of a spike train.
In this paper, we rigorously explore from first principles the transition and modeling from the standard biophysical picture of a single action potential to its representation as a spike in a spike train. In particular, we adopt a functional-analytic framework, using Schwartz distribution theory to represent spike trains as generalized Dirac delta functions acting on smooth test functions. We then show how and why this representation transcends a purely descriptive formalism to support deep downstream analysis and modeling of spike train neural dynamics in a mathematically consistent way.
The ideas and content in this paper draw on several decades of work in theoretical and computational neuroscience and neural dynamics. Representative treatments ranging from early point-process formulations of spike trains to modern dynamical systems and network models can be found in [DayanAbbott2001], [GerstnerKistler2002], [GerstnerKistlerNaudPaninski2014], [Tuckwell1988], [PerkelGersteinMoore1967], [Knight1972], [BrunelHakim1999], and [ErmentroutTerman2010], among many others. See also the references within these cited works.
The primary contribution of the present work is not the introduction of new neurophysiological mechanisms or descriptions, but a first-principles reconstruction of these concepts within a unified framework grounded in functional analysis and distribution theory [botelho2025, brayman2024, aksoy2024, muscat2024, clason2020]. By making the underlying mathematical structure explicit, the paper clarifies the sense in which spike trains are treated as distributions: mathematical objects defined not as functions or by their height at a single point, but by their interactions as unique events with smooth test functions that act as ’probes’. This framework provides a rigorous basis for treating spikes as mathematical objects that isolate timing and weight as the primary variables driving neural dynamics while maintaining consistency with neurophysiologically meaningful dynamical systems.
The intent is to provide a self-contained, technically precise reference that bridges what is often a communication barrier between the physiological and mathematical perspectives. It offers readers a rigorous foundation for engaging with and extending the extensive literature on spike-based neural dynamics.
2 The Representational Objective
Assume neuron $i$ produces a sequence of spike times given by
$0<t_{i}^{1}<t_{i}^{2}<\cdots<t_{i}^{k}\cdots t_{i}^{K}$
This collection, or set, of spike times for neuron $i$ is what we refer to as its spike train. (A raster plot is a matrix of spike trains, i.e., spike trains for different neurons superimposed on the same time axis.) A spike train is not naturally a classical function of the form
$f:\mathbb{R}\rightarrow\mathbb{R},\text{ }t\mapsto\mathbb{R}$
for two reasons. First, the events are conceptual idealizations of instantaneous events and have zero duration. Second, and critically, what matters operationally is not the value or amplitude of the spike at an individual time $t$, but how spikes contribute to or influence downstream quantities, for example, synaptic currents, counts, a kernel, a low-pass filter, or averages over a time window. The effects of spikes on these classes of functions are necessarily almost always integrals or sums.
To define a spike train, we need to define an object $s_{i}$ such that any physically meaningful operation we do with it can be expressed as an integral or sum paired with a test function $\varphi(t)$ that measures the effect of $s_{i}$ on the operation. A valid test function will need to satisfy a number of formal mathematical properties and will be of practical neurophysiological utility, as in the examples introduced in the previous paragraph.
Formally, we want to understand, or compute, the degree to which a spike at $t_{i}^{k}$ contributes to or affects $\varphi(t_{i}^{k})$. For multiple spikes, we want the contributions to add, such that
$\quad\langle s_{i},\varphi\rangle:=\sum_{k}\varphi(t_{i}^{k})$ (2.1)
Equation 2.1 is then the operational requirement for a spike train. A set $s_{i}$ paired with a test function $\varphi(t)$ such that $s_{i}$ is defined by how it samples and evaluates the summed effect of $\varphi(t)$ only at specific times.
3 Why the Dirac delta is not a classical (Lebesgue) function
We will make use of the Dirac delta as the unique mathematical object that formally encodes the spike train $s_{i}$ to sample a test function $\varphi(t_{i}^{k})$ in such a way that it satisfies the requirement defined by Equation 2.1. To understand how the Dirac delta does this, and why the mathematics remains consistent, one has to first deeply understand what the Dirac delta is, because it is not a function in the usual way we think of functions as maps.
The way the Dirac delta is typically introduced in a neurophysiological context to model spikes can be misleading. So we will first formally define the Dirac delta from first principles and review some of its properties, then return to its use to encode spike trains.
The Dirac delta $\delta$ is often introduced by associating it with three key properties. Namely:
$\delta(t-a)=0\text{ for }t\neq a$ The delta function is zero everywhere except $t=a\displaystyle\text${The delta function is zero everywhere except at }t=a (3.1a)
$\delta(a)=\infty$ Its magnitude is infinitely large $t=a\displaystyle\text${Its magnitude is infinitely large at }t=a (3.1c)
$\int_{-\infty}^{\infty}\delta(t-a)\,dt=1$ It has an area equal to unity (3.1e)
In other words, it is a rectangle that is infinitely high and infinitely narrow exactly and only at $t=a$.
Almost immediately, one gets a sense that this is not a function in the ordinary sense. By an ‘ordinary’ or classical function, here, we mean an integrable Lebesgue function. A function that assigns a real number to (almost) every time $t$, in a way that is well-defined pointwise and compatible with Lebesgue integration. Formally, this means that $f:\mathbb{R}\to\mathbb{R}$ such that for each $t\in\mathbb{R}$ (or for almost every $t$, not including any discontinuities), the value $f(t)$ exists as a real number. This type of function is measurable with respect to the Lebesgue measure, so the integration of the function $\int f(t)\,dt$ makes sense and can be defined. (There are other conditions and implications, but for our purposes here, this is sufficient to make our point.)
If we try to treat $\delta$ as a Lebesgue function, we run into an issue. If $\delta=0$ for all $t\neq a$, then $\delta$ differs from the zero function $g(t)=0$ $\forall t$, i.e., the function that equals zero for all inputs, only at a single point. But changing a function on a single point does not change its Lebesgue integral. This means that
$\int\delta(t-a)\,dt=0,$
which contradicts the area =1=1 claim. There is no classical function with those properties under ordinary integration.
The Lebesgue integral does not care about what happens at individual points. What matters is what happens on sets with nonzero measure, such as length, area, volume, etc. So for example, the interval $[0,1]$ has a Lebesgue measure $\mu([0,1])=1$. An interval $[a,b]$ has $\mu([a,b])=b-a$. But any finite countable set of points, for example, $\{3,7,9,56\}$ or $\{a\}$ or $\{0\}$ itself, has a measure equal to zero, i.e., $\mu(\{a\})=0$.
In the context of Lebesgue integration and how it is formally defined, any two functions that differ at a single point, including $\delta$ and $g(t)=0$ $\forall t$, have the same integral. To see this, consider an arbitrary function
$f(t)=\begin{cases}M&t=a\\ 0&t\neq a\end{cases}$
for $M\in\mathbb{R}$. $f(t)$ and $g(t)$ differ only at the single point $t=a$, and both have measures $\mu(f)=\mu(g)=0$. This implies that
$\int f(t)dt=\int g(t)dt$
But we know that the integral of the zero function is zero: $\int g(t)dt=0$, therefore $\int f(t)dt=0$.
This holds regardless of the value of $M$. If $M\rightarrow\infty$, then $\mu((f(M))=0$. In other words, values on sets of measure zero do not contribute to the integral. This contradicts the property that $\int\delta(t-a)dt=1$ for any value of $(t-a)\in\mathbb{R}$.
The way to get around this issue is by satisfying the operational requirement of Equation 2.1. We will do this by formally defining the Dirac delta $\delta$ as an object called a Schwartz distribution.
Although the Dirac delta is not an ordinary Lebesgue function and cannot be treated as one, it can be obtained as a limit of ordinary functions in the sense of Schwartz distributions. In other words, it can be approximated arbitrarily well by ordinary functions, provided the limit is understood in the context of distributions. This is the basis for the intuitive notion often used to introduce the Dirac delta. But to do this properly, we first need to understand what Schwartz distributions are. So we will hold off on showing this until Section 6 below.
4 Schwartz distributions
Schwartz distributions were introduced and formalized by Laurent Schwartz (1915-2002) in the late 1940s as a new class of mathematical objects that extended functions in order to allow differentiation and Fourier analysis for singular objects like the Dirac delta [botelho2025], [brayman2024], [aksoy2024], [muscat2024], [clason2020]. They are distinct from, and not to be confused with, statistical distributions, which most neuroscientists are much more familiar with.
Formally, a Schwartz distribution $T$ is a continuous linear functional on a space of test functions. We will carefully examine each part of this definition.
4.1 Function vs functional
In the most straight forward sense, a function is a map from numbers to numbers:
$f:\mathbb{R}\to\mathbb{R},\quad x\mapsto f(x)$
(More accurately a function maps elements in a domain set to a codomain set. They can be numbers but many other things too.) A functional maps functions to numbers:
$F:\mathcal{F}\to\mathbb{R},\quad\varphi\mapsto F(\varphi)$
where $\varphi$ represents a function. In other words, the input to a functional is itself a function.
A well-known and easy example of a functional is the definite integral. It takes a function as its input argument and returns the area under the curve defined by the function over the specified limits. For example
$F(\varphi):=\int_{-\infty}^{\infty}\varphi(t)\,dt$ integrated across the entire number line (4.1a)
$F(\varphi):=\int_{a}^{b}\varphi(t)\,dt$ $\text{integrated on the closed intervarl }[a,b]$ (4.1b)
By definition, the Dirac delta is defined as the functional that evaluates a function at a point.
$F(\varphi)=\langle\delta(t-a),\varphi\rangle=\delta_{a}(\varphi):=\varphi(a)$ (4.2)
The Dirac delta is a mathematical object (a functional) that takes a test function $\varphi(t)$ and returns a number, explicitly, the value of $\varphi(t)$ at the point $a$. Notice how the pairing $\langle\delta(t-a),\varphi\rangle$ produces a number, the value of the test function at $t=a$, and satisfies the operational requirement set up in Equation 2.1. This define the Dirac delta for the single-valued case, but not (yet) a spike train $s_{i}$, which we work up to below.
The transition from functions to functionals reflects a fundamental shift in how a mathematical object is conceptualized. A classical function maps a unique output for every input. A functional, however, provides a measurement for every test. In this framework, the test function $\varphi(t)$ acts as a mathematical probe. There is no interaction with a distribution directly. A distribution is is characterized only by observing how it ”weights” or filters different probes (test functions) that interact with it.
4.2 Linearitiy of functionals
Functionals also satisfy linearity, in the usual sense we think of linearity in vector spaces. Linearity must satisfy
$F(\alpha\varphi+\beta\psi)=\alpha F(\varphi)+\beta F(\psi)\quad\text{for all }\alpha,\beta\in\mathbb{R}$
An integral $F(\varphi)=\int\varphi(t)\,dt$ is linear because it can be written as
$F(\alpha\varphi+\beta\psi)=\int(\alpha\varphi+\beta\psi)\,dt=\alpha\int\varphi\,dt+\beta\int\psi\,dt=\alpha F(\varphi)+\beta F(\psi)$
Similarly, the Direct delta is linear because it satisfies
$\delta_{a}(\alpha\varphi+\beta\psi)=(\alpha\varphi+\beta\psi)(a)=\alpha\varphi(a)+\beta\psi(a)=\alpha\delta_{a}(\varphi)+\beta\delta_{a}(\psi)$
From the perspective of modeling spike trains using the Dirac delta, linearity is critical because it is the property that supports the superposition of spike contributions and the resultant effect on the test function under consideration.
4.3 The space of test functions: Definition of support
Test functions are smooth, well-behaved functions that probe the distribution. We use test functions to ask questions and observe the effects of the distribution on them. Schwartz distributions, like the Dirac delta, are not defined pointwise. They do not evaluate an input that maps to an output like a regular function (see the discussion above in Section 3.) They only make sense in context with how they interact with test functions.
Test functions $\varphi(t)$ are not biophysically measurable or derived variables like the membrane potential, synaptic currents, or ionic conductances. They are any smooth time-localized filter that can be properly defined that can be used to interrogate the distribution. Explicitly, in our case, the distribution being constructed is a spike train encoded by Dirac deltas. We need to restrict the space of test functions to only those with these properties because, for arbitrary functions, evaluation at a point may not be defined, differentiation may be undefined, and convergence may not hold. All of which we need in the interaction between a functional and a test function. Below, we rigorously formalize these requirements.
To model a spike train, we can restrict the space of test functions to $C_{c}^{\infty}$, the space of smooth functions with compact support on $\mathbb{R}$. $C$ implies continuous functions, while $C^{\infty}$ implies continuous functions for which all derivatives exist. In other words, they are infinitely differentiable. This makes the space of test functions not just continuous but smooth, a stronger condition than continuity; i.e., all smooth functions are continuous, but not all continuous functions are smooth. Formally,
$f\in C^{\infty}(\mathbb{R})\quad\Longleftrightarrow\quad\frac{d^{k}f}{dt^{k}}(t)\ \text{exists for all }k\in\mathbb{N}$
The subscript in $C_{c}^{\infty}$ means the set of smooth functions that are compactly supported, meaning if $\varphi(t)\in C_{c}^{\infty}\rightarrow\varphi(t)=0$ outside some finite interval. Understanding what the support of a function is, and how it differs between ordinary (Lebesgue) functions and Schwartz distributions, is not just a technical mathematical detail; it is at the core of understanding why the Dirac delta works as a model of spike trains.
The requirement for test functions to be smooth ($C^{\infty}$) is a strategic choice that shifts the burden of differentiation. Because many distributions (like the spike train $s_{i}$) are singular events and lack classical derivatives, distribution theory transfers the operation of differentiation onto the test function. Because $\varphi(t)$is infinitely differentiable, the definition of differentiation of the spike train is dependent on the derivatives of the test function. This ensures that the mathematical machinery never breaks down when encountering singularity associated with the events that make up a distributional like the spike train. We discuss this in some detail in Section 7.4 at the end of the paper.
We will first define the support of ordinary functions, which applies to test functions $\varphi(t)$. We will then show why that definition fails for distributions, and finally, how we can adjust the definition of support to accommodate distributions. What will come out of this exercise is an understanding of what it means for Dirac deltas to represent and encode spikes in a spike train, and why the concept of Schwartz distributions only makes sense in context with their effects on test functions.
Let $f:\mathbb{R}\to\mathbb{R}$ represent a classical Lebesgue function (see Section 3). We define the support of such a function $\mathrm{supp}(f)$, as the closure of the subset of the domain on which $f$ is nonzero:
$\mathrm{supp}(f):=\overline{\{t\in\mathbb{R}:f(t)\neq 0\}}$ (4.3)
There is a subtle yet critical notion here that can easily be missed. This definition relies on pointwise values. The support is defined as a subset of the function’s domain (the subset where it is nonzero), and includes boundary points wherever the function is nonzero arbitrarily close to them. This means that support is not defined by where a function is nonzero at that exact point. It is defined by where it is nonzero arbitrarily close to that point. Or put another way, a point in the domain of $f(t)$ belongs to the support if the function cannot be made identically zero in any arbitrarily small neighborhood around that point. This definition is designed to capture the notion of local influence, not pointwise values of $f(t)$ for some value of $t$. If this was not the case, and we attempted to define support on a pointwise basis for all nonzero values, i.e., $\{t:f(t)\neq 0\}$, then boudary points would be arbitrarily excluded, since any nonzero value of $f(t)$ could be made to approach the boundary point as close as we like. There would be no way to non-arbitrarily define where the function goes from being nonzero to zero. This would produce downstream inconsistencies. For example, integration by parts would fail. And, most importantly for us in this paper, the definition of support would not generalize to Schwartz distributions.
For example, if $f(t)=e^{-t^{2}}$, then $f(t)\neq 0$ for all $t$, so $\mathrm{supp}(f)=\mathbb{R}$. Given some other function $f(t)$ that is zero outside the closed interval $[a,b]$, its support is $\mathrm{supp}(f)=[a,b]$.
If $f(t)\neq 0$ on the open interval $(0,1)$, and $f(0)=f(1)=0$, the support would still be the closed interval $[0,1]$, i.e., $\mathrm{supp}(f)=[a,b]$. The nonzero set in this case is $\{t:f(t)\neq 0\}=(0,1)$, but the closure of the open set is the closed set $[0,1]$ because every neighborhood of 0 contains points $t>0$ (to the right of 0) even though $f(0)=0$. And every neighborhood of $1$ contains points $t<1$ (to the left of 1) even though $f(1)=0$. So the points 0 and $1$ are limit points of the nonzero open set $(0,1)$, and are therefore included in the support of the function.
But if $f(t)\neq 0$ for $f(0)$ and $f(1)$, then it is nonzero only at isolated points, and differs from the the zero function $g(t)0$ $\forall t\in\mathbb{R}$ on a finite set (like we described above). Such a function has a Lebesgue support equal to the empty set. This is why attempting to use the Lebesgue definition of support fails for the Dirac delta and, more generally, for Schwartz distributions. Instead, the support of a distribution is defined operationally by how it acts on test functions.
The support of a distribution $T$, $\mathrm{supp}(T)$, is defined as the complement of the largest open set on which $T$ vanishes. Formally, a point $t_{0}\in\mathbb{R}$ is not in $\mathrm{supp}(T)$ if and only if there exists an open neighborhood $U\ni t_{0}$ such that
$\langle T,\varphi\rangle=0\quad\text{for all test functions }\varphi\text{ with }\mathrm{supp}(\varphi)\subset U.$
[…]
We can now use this definition to derive the support of the Dirac delta $\delta(t-a)$. Assume that $\delta(t-a)=1$ and zero otherwise. First, we show how $\delta$ vanishes away from $a$. Take any point $t_{0}\neq a$. There exists an open interval $U$ containing $t_{0}$ but not containing $a$. Next, assume any test function $\varphi$ with $\mathrm{supp}(\varphi)\subset U$. Since $a\notin\mathrm{supp}(\varphi)$, we have $\varphi(a)=0$. By definition, though,
$\langle\delta(t-a),\varphi\rangle=\varphi(a)=0$
This implies that there is an open neighborhood around every point $t_{0}\neq a$ on which $\delta(t-a)$ acts as zero. Therefore, $t_{0}\notin\mathrm{supp}(\delta(t-a))$ for all $t_{0}\neq a$.
In contrast, we can prove that $\delta$ does not vanish at $a$. This time, consider $t_{0}=a$. Take any open neighborhood $U\ni a$. We assume there exists a test function $\varphi\in C_{c}^{\infty}$ such that $\mathrm{supp}(\varphi)\subset U$, and $\varphi(a)\neq 0$. This means that $\langle\delta(t-a),\varphi\rangle=\varphi(a)\neq 0$. There is no neighborhood of $a$ on which $\delta(t-a)$ vanishes, so it must be that $a\in\mathrm{supp}(\delta(t-a))$.
Combining the two conditions then implies that $\mathrm{supp}(\delta(t-a))=\{a\}$. $\delta(t-a)$ is completely insensitive to anything that happens outside arbitrarily small neighborhoods of time $a$.
Putting all the pieces together, we represent a spike train $s_{i}$ as the distribution
$s_{i}(t)=\sum_{k}\delta(t-t_{i}^{k}),$ (4.4)
where the sum is understood in the sense of distributions. Since each Dirac delta $\delta(t-t_{i}^{k})$ has support $\{t_{i}^{k}\}$ and the spike times are locally finite, the support of the spike train is given by
$\mathrm{supp}(s_{i})=\overline{\{t_{i}^{k}:k\in\mathbb{N}\}}.$ (4.5)
This restricts support of the spike train to only the spike times. In particular, $s_{i}$ is not a function of time in the classical sense, but a distribution whose action on test functions samples them at the discrete event times $\{t_{i}^{k}\}$.
4.4 The space of test functions: Continuity of functionals
Recall from the last section that because the space of test functions $C_{c}^{\infty}$ are infinitely differentiable and smooth, it immediately implies continuity.
But from the perspective of functionals, continuity means that if a sequence of test functions converges smoothly and locally to a limit, then the outputs produced by the functional also converge. In contrast to the usual function continuity, this is a different meaning and definition of continuity. This is necessary to ensure stability and provide physical meaning and interpretation to modeled variables and systems.
Given a set of test functions $\{\varphi_{1},\varphi_{2},\cdots\varphi_{n}\cdots\varphi_{N}\}\in C_{c}^{\infty}(\mathbb{R)}$, the set of functions converge to a limit function $\varphi$,
$\varphi_{n}\to\varphi\quad\Longrightarrow\quad F(\varphi_{n})\to F(\varphi)$
if the functions converge uniformly:
$\sup_{t}|\varphi_{n}(t)-\varphi(t)|\to 0$
all their derivatives converge uniformly:
$\sup_{t}|D^{k}(\varphi_{n}-\varphi)(t)|\to 0\quad\text{for all }k$
and the supports remain bounded in a compact set, i.e., functions do not explode to infinity. If these conditions are met, it ensures that any $\varphi_{n}$ is indistinguishable from $\varphi$ even under arbitrarily fine local inspection. Note that ’sup’ here is not ’support’ but ’supremum’, meaning least upper bound, in this case taken over the entire variable $t$, which is why we write $\sup_{t}$. Notaionally, we are also using $D^{k}\equiv\frac{d^{k}\varphi}{dt^{k}}$. So writing $\sup_{t}|D^{k}(\varphi_{n}-\varphi)(t)|$ means the largest value ($\equiv$ the smallest possible upper bound) of all the derivatives of $(\varphi_{n}-\varphi)(t)$ as $t$ ranges over all real numbers, i.e., $t\in\mathbb{R}$.
In the case of the Dirac delta defined by Equation 4.2, where $F(\varphi):=\delta_{a}(\varphi):=\varphi(a)$, if $\varphi_{n}\rightarrow\varphi$, then $\varphi_{n}(a)\to\varphi(a)$ for any value $t=a$. And
$F(\varphi_{n})=\varphi_{n}(a)\to\varphi(a)=F(\varphi)$
implying that the Dirac delta is, as required, a continous functional.
5 Integral notation for the Dirac delta
As a notational comment, $\delta$ is often written in integral notation to express the operational definition given in Equation 4.2:
$\int_{-\infty}^{\infty}\varphi(t)\,\delta(t-a)\,dt=\varphi(a):=\delta_{a}(\varphi)$ (5.1)
When containing a functional, this expression should not be confused with the normal integration between two (Lebesgue) functions, where we multiply the two functions point by point and add up the products over time. In the context of functionals it simply means applying a linear functional, which for the Dirac delta is defined to return $\varphi(a)$.
While seemingly unnecessary, this notation has a number of important consequential advantages. It allows us to deal with smooth functions and spikes in a single functional framework, and maintains consistency when working with convolutions and integration by parts. We show this below.
6 How the Dirac Delta Arises as a Limit of Ordinary Functions
In this section, we return to the notion introduced in Section 3 of the Dirac delta representing an infinitely narrow, infinitely tall rectangle with a unitary area. Mathematically, this results from taking the limit of an ordinary function, but we need to use the mathematical machinery of distributions and test functions to do so properly.
We beginning by letting $\rho(t)\geq 0$ be a smooth function with:
$\int_{-\infty}^{\infty}\rho(t)\,dt=1.$
We then define a scaled family of related functions by
$\rho_{\varepsilon}(t):=\frac{1}{\varepsilon}\rho\!\left(\frac{t}{\varepsilon}\right).$
for the parameter $\epsilon$.
Notice how the area reamins equal to 1, since
$\int\rho_{\varepsilon}(t)\,dt=\int\frac{1}{\varepsilon}\rho\!\left(\frac{t}{\varepsilon}\right)dt=\int\rho(u)\,du=1\text{ where, }u=\frac{t}{\epsilon}$
The functions progressively cluster around $t=0$ as $\varepsilon\to 0$. To center the functions at the point $a$, we define $\rho_{\varepsilon}(t-a)$.
For any test function $\varphi\in C_{c}^{\infty}$, we can write
$\int\varphi(t)\rho_{\varepsilon}(t-a)\,dt.$
and substituting for a shifted center point $t=a+\varepsilon u$, this becomes
$\int\varphi(a+\varepsilon u)\rho(u)\,du.$
As $\varepsilon\to 0$, $\varphi(a+\varepsilon u)\to\varphi(a)$ pointwise, and because $\varphi$ is smooth and compactly supported (see Sections 4.3 and 4.4), convergence applies, giving
$\lim_{\varepsilon\to 0}\int\varphi(t)\rho_{\varepsilon}(t-a)\,dt=\int\varphi(a)\rho(u)\,du=\varphi(a)\int\rho(u)\,du=\varphi(a).$
So $\rho_{\varepsilon}(t-a)$ converges to $\delta_{a}$ as distributions:
$\rho_{\varepsilon}(t-a)\xrightarrow{\ \mathcal{D}^{\prime}\ }\delta_{a}$ (6.1)
(A necessary technical note: Following standard functional analysis notation, we write $\mathcal{D}^{\prime}$ to indicate the space of (Schwartz) distributinos. The space of distributions $\mathcal{D^{\prime}}(\mathbb{R})$ is the dual space to the space of test functions: $\mathcal{D}(\mathbb{R})=C_{c}^{\infty}(\mathbb{R})$. $C_{c}^{\infty}(\mathbb{R})$ is formally a vector space, and the dual space of distributions is the space of linear functionals. So writing $\rho_{\varepsilon}(t-a)\xrightarrow{\ \mathcal{D}^{\prime}\ }\delta_{a}$ in Equation 6.1 implies that in the limit the functions converge in the sense of distributions.)
This is the meaning behind the idea of a Dirac delta as an infinitely tall, infinitely narrow, rectangle with area equal to one.
7 Putting it all Together: The Dynamics of Spike Trains
In this last section, we apply the framework developed above to the dynamics of spike trains. Fundmentally, the purpose of this formalism is not just to represent spikes more accurately, but to enable mathematically rigorous and consistent modeling and analysis of neuronal dynamics that cannot be treated coherently without it.
7.1 Linear summation of spike distributions
Recall from Equation 4.2 that $\langle\delta(t-a),\varphi\rangle=\delta_{a}(\varphi):=\varphi(a)$ Given that distributions are linear functionals (Section 4.2), if $T_{1}$ and $T_{2}$ are distributions, then $T_{1}+T_{2}$ is defined by
$\langle T_{1}+T_{2},\varphi\rangle=\langle T_{1},\varphi\rangle+\langle T_{2},\varphi\rangle$ (7.1)
c.f. also Equation 2.1. In Equation 4.4 we defined a spike train as
$s_{i}(t)=\sum_{k}\delta(t-t_{i}^{k})$ Equation 4.4 (7.2a)
This is not a function, it is a sum of linear functionals. Due to linearity, the action of this sum on a test function $\varphi\in C_{c}^{\infty}$ is
$\left\langle\sum_{k}\delta(t-t_{i}^{k}),\varphi\right\rangle=\sum_{k}\langle\delta(t-t_{i}^{k}),\varphi\rangle=\sum_{k}\varphi(t_{i}^{k})$ (7.3)
Equation 7.3 is interpreted as sampling the test function at every spike and adding up the results, which is the intended behavior.
Whereas, Equation 4.2 satisfies the operational requirment we set up in Equation 2.1 for the single value case, Equation 7.3 equivalently does so for a linear combination of functionals.
7.2 Local finiteness and well-defined spike trains
We address a technical consideration to ensure that the spike train representation is mathematically well-defined. Because distributions are defined by their action on compactly supported test functions, the only way an infinite sum of spike events could fail to be well-defined is if infinitely many spikes contribute within a finite time window. This condition is avoided by assuming local finiteness, that is, only finitely many spikes may occur in any bounded interval of time, even though the spike train may contain infinitely many events overall, athematically speaking at least. In practice, of course, any spike train is necessarily finite. But this requirment ensures mathematical consistency.
Intentionally, test functions $\varphi\in C_{c}^{\infty}(\mathbb{R})$ have compact support (c.f. Section s 4.3 and 4.4). This ensures that there exists a finite interval $[a,b]$ such that $\varphi(t)=0$ outside $[a,b]$. Because $t\in\mathbb{R}$, the support of $\varphi$ takes the infinite sum
$\sum_{k}^{\infty}\varphi(t_{i}^{k})$
and restricts it to
$\sum_{k:\,t_{i}^{k}\in[a,b]}\varphi(t_{i}^{k})$
since every term $t_{i}^{k}\notin[a,b]$ is zero. The only remaining condition we need to impose is that that there not be infinitely many spike times inside $[a,b]$. So we assume local fititeness such that
$\{k:t_{i}^{k}\in[a,b]\}<\infty\quad\text{for every bounded interval }[a,b]$
This guarantees that for every test function only finite many nonzero terms (spikes) appear in the sum, making the sum finite and precise. Otherwise, the sum becomes a conditionally convergent (or divergent) infinite series due to Riemann rearrangement, whereby any conditionally convergent infinite series of real numbers can be rearranged to converge to any value, or to diverge. From a neurophysiological and neural dynamics perspective, this constraint is, of course, reasonable, since we would never record an infinite number of spikes in a finite time window.
7.3 Convolution and spike train dynamics
In this section, we show how the spike-train representation naturally gives rise to the standard convolution dynamical description of synaptic and membrane processes. Once spikes are modeled as distributions, i.e., instantaneous events defined only through their action on test functions, dynamics is no longer a pointwise evaluation in time, but linear filtering via convolution. Each spike produces a time-shifted copy of a fixed response kernel, and multiple spikes then combine by superposition. This formalism naturally explains why synaptic dynamics take the form of shifted impulse responses, and provides the mathematical machinery to study it. We will use this example of synaptic dynamics to develop the intuition.
We want to construct a spike train $s_{i}$ such that each spike at time $t_{i}^{k}$ produces a stereotyped postsynaptic response waveform kernel $K$ that gets triggered at that time. Consistent with the formalism we laid out for Equation 2.1, the subscript $i$ indexes individual neurons, while the superscript $k$ indexes individual spike events within the spike train of neuron $i$. Leveraging the linear summation of multiple occuring spiking events, we can write an expression for synaptic current $I_{syn}$ as
$I_{\text{syn}}(t)=\sum_{k}K(t-t_{i}^{k})$ (7.4)
Neurophysiologically, $K$ here represents the elemetary postsynaptic current contributed by each presynaptic neuron on the postsynaptic neuron $i$. $I_{syn}(t)$ is the linear supperposition of individual currents, i.e., the total synaptic current. For example, a standard form for a postsynaptic current kernel for ionotropic $Cl^{-}$ gamma-aminobutyric acid (GABA) and $Na^{+}$ $\alpha$-amino-3-hydroxy-5-methyl-4-isoxazolepropionic acid (AMPA) ion channels is
$K(t)=g_{\max}\,\bigl(e^{-t/\tau_{d}}-e^{-t/\tau_{r}}\bigr)H(t)$
where $\tau_{r}$ and $\tau_{d}$ are the waveform rise and decay time constants, $g_{max}$ is the maximum or peak channel conductance, and $H(t)$ is the Heaviside step function which acts as a switch [dayan2001theoretical], [destexhe1994efficient], [destexhe1998kinetic].
By modeling the spike train as a set of Dirac deltas $s_{i}(t)=\sum_{k}\delta(t-t_{i}^{k})$, the convolution of the kernel with the delta spike train representation computes to our expression of $I_{syn}$ in Equation 7.4.
To see this, we define convolution for an ordinary function $f$ with a kernel $K$ in the usual way as
$(K*f)(t):=\int_{-\infty}^{\infty}K(t-\tau)\,f(\tau)\,d\tau$ (7.5)
For a fixed (current) time $t$, we take into account all past input times $\tau$. For each $f(\tau)$ we weight it by how much it contributes to the output at the computed time $t$ using the kernel $K(t-\tau)$. Finally, we sum (integrate) over all time points $\tau$. This is a linear time invariant filter.
For a single spike at $\tau=a$ represented as a Dirac delta, $f(\tau)=\delta(\tau-a)$, substituting into Equation 7.5 yields
$(K*\delta(\tau-a))(t)=\int_{-\infty}^{\infty}K(t-\tau)\,\delta(\tau-a)\,d\tau.$ (7.6a)
Recall from Equation 5.1 in Section 5 that
$\int_{-\infty}^{\infty}\varphi(\tau)\,\delta(\tau-a)\,d\tau=\varphi(a)$
But here we have an explicit form for $\varphi(t):=K(t-\tau)$ so putting this back into Equation 7.6a gives
$(K*\delta(\tau-a))(t)=\int_{-\infty}^{\infty}K(t-\tau)\,\delta(\tau-a)\,d\tau=K(t-a)$ (7.6b)
A spike at $\tau=a$ results in a postsynaptic response shifted to start at $\tau=a$.
Extending this to a series of spikes, i.e., a spike train $s_{i}(\tau)=\sum_{k}\delta(\tau-t_{i}^{k})$, assuming linearitiy (Sections 4.2 and 7.1) and local finiteness (Section 7.2), we compute the convolution in the same way (dropping the integration limits for notational simplicity):
$(K*s_{i})(t)=\int K(t-\tau)\,s_{i}(\tau)\,d\tau=\int K(t-\tau)\left(\sum_{k}\delta(\tau-t_{i}^{k})\right)d\tau$ (7.7)
Notice how each term in Equation 7.7 is
$\int K(t-\tau)\,\delta(\tau-t_{i}^{k})\,d\tau=K(t-t_{i}^{k})$
which is the multi-spike version of Equation 7.6b we derived above. This implies that
$(K*s_{i})(t)=\sum_{k}K(t-t_{i}^{k})$ (7.8)
Equation 7.8 is the explicit interpretation of convolution in the context of spike train dynamics. Each presynaptic spike at time $t_{i}^{k}$ contributes a time-shifted copy of the elementary postsynaptic response $K$, and the total synaptic current at time $t$ is obtained by summing these contributions. Importantly, the process does not rely on treating spikes as functions of time; it follows directly from representing spikes as distributions and applying linear filtering. Convolution emerges as the unique operation that maps discrete spike events to continuous synaptic dynamics that preserves linearity, causality, and temporal locality.
In the context of convolution, the Dirac delta acts as a form of time translation. When a synaptic kernel $K$ is convolved with a time-shifted delta $\delta(t-a)$, the delta essentially translates the entire shape of the kernel along the timeline to exactly $t=a$. This functionally transforms a complex integral into a simple copy-paste operation: for every spike time in the collection of events, the distributiional operation pastes a copy of the synaptic response into the total current, allowing discrete events to be encoded into continuous-time dynamics.
This formalism shows why convolution provides the natural dynamical description for event-driven spiking neuronal processes. Once spikes are modeled as instantaneous events in the sense of distributions, linear filtering by convolution is the only operation able to encode discrete spike times into continuous-time responses. Synaptic currents, membrane potentials, and other downstream variables inherit their temporal structure directly from the kernel $K$ and the geometry of spike timing.
7.4 Distributional differentiation of spike trains
We end the paper by showing another example that emphasizes the unique gains of a distribution theory formalism to neuronal spike trains. Although a spike train has no pointwise derivative, its derivative is well-defined in the distributional sense. By definition, the derivative $s_{i}^{\prime}$ acts on a test function $\varphi$ as $\langle s_{i}^{\prime},\varphi\rangle=-\langle s_{i},\varphi^{\prime}\rangle$, producing $\sum_{k}\varphi^{\prime}(t_{i}^{k})$. In other words, differentiation of a spike train does not measure changes in the spikes themselves, which has no meaning, but rather local changes in the test function at the event times. We will develop this intuition carefully.
A spike train is a set of instantaenous events, so what does it even mean to differentiate a set like that? When we differentiate ordinary functions, derivatives have phyical and geometric interpretations, such as the rate of change, or slope. But distributions like the spike train are not ordinary functions. They have no meaning or values between events, or concepts like slope. This necessitates a change in perspective that ultimately results in a powerful set of mathematical ideas and tools to analyze and understand (in our case) very real world physical and physiological properties of neural signaling and dynamics.
This shift in perspective involves asking not how do we differentitae the distribution or what that might mean. The spike train cannot be directly differentiated. Instead, we differentiate the test functions and ask how does the spike train respond.
From the operational requirement set in Equation 2.1,
$\langle s_{i},\varphi\rangle=\sum_{k}\varphi(t_{i}^{k})$
with $s_{i}(t)=\sum_{k}\delta(t-t_{i}^{k})$ (Equation 4.4). The spike train samples the test function at the event times.
For any distribution $T$, its derivative $T^{\prime}$ is defined as
$\langle T^{\prime},\varphi\rangle:=-\langle T,\varphi^{\prime}\rangle$ (7.9)
This definition is consitent with the ordinary notion of differentiation if $T$ were a smooth function, and preserves integration by parts, but avoids pointwise evaluation. We look at this more deeply (and why the negative sign in the definition needs to be there) below.
For a spike train, $T=s_{i}$
$\langle s_{i}^{\prime},\varphi\rangle$ $=-\langle s_{i},\varphi^{\prime}\rangle$ (7.10a)
$\langle s_{i}^{\prime},\varphi\rangle$ $=-\sum_{k}\varphi^{\prime}(t_{i}^{k})$ (7.10b)
The intepretation of Equation 7.10 is that the original spike train $s_{i}$ samples values of the test function $\varphi(t)$ at the event times, while the derivative of the spike train, $s_{i}^{\prime}$, samples values of the derivative of the test function $\varphi^{\prime}(t)$. In other words, $s_{i}^{\prime}$ samples values of the slope of $\varphi(t)$. This means that the derivatives of the spike train (by definitiion) measures how the test function changes at the event times.
[…]
Two closing technical comments: First, if $\varphi(t)$ is not differentiable, then $\varphi^{\prime}(t)$ would be undefined, making $s_{i}^{\prime}$ undefined (directly due to its definition). So this operation necessitates all the work we did above to justify why $\varphi(t)\in C_{c}^{\infty}$, and as a result needs to be smooth (in a mathematically formal sense).
Second, the minus sign in Equation 7.9 results directly from integration by parts. It is worth cosidering why it arrises in some detail, because it provides a deep intuative insight.
Assume some function $f(t)\in C^{\infty}$ is an ordinary smooth function that is smooth but not necessarily nonzero everywhere and possibly grows to infinity. We still require a test function $\varphi(t)\in C_{c}^{\infty}$, i.e., smooth and compactly supported. We can choose to view $f(t)$ as a distribution by defining
$T_{f}(\varphi):=\int f(t)\,\varphi(t)\,dt$ (7.11)
Then, taking Equation 7.9 into consideration, the distributional derivative of $T_{f}(\varphi)$ is
$\langle T_{f}^{\prime},\varphi\rangle:=-\langle T_{f},\varphi^{\prime}\rangle=-\int f(t)\varphi^{\prime}(t)\,dt$ (7.12)
Because $f(t)$ is an ordinary smooth function, and $\varphi(t)$ is smooth and compactly supported, we can integrate by parts. First, notice that due to the product rule, the derivative of the product of $f(t)\varphi(t)$ is
$\frac{d}{dt}\bigl(f(t)\varphi(t)\bigr)=f^{\prime}(t)\varphi(t)+f(t)\varphi^{\prime}(t)$
Integrating both sides of this expression over $\mathbb{R}$ produces
$\int_{-\infty}^{\infty}\frac{d}{dt}\bigl(f(t)\varphi(t)\bigr)\,dt=\int_{-\infty}^{\infty}\Bigl(f^{\prime}(t)\varphi(t)+f(t)\varphi^{\prime}(t)\Bigr)\,dt$ (7.13a)
If we consider the left hand side of Equation 7.13a, the fundamental theorem of calculus results in an evaluation of the boundary conditions
$\int_{-\infty}^{\infty}\frac{d}{dt}\bigl(f(t)\varphi(t)\bigr)\,dt=\Bigl[f(t)\varphi(t)\Bigr]_{t=-\infty}^{t=\infty}$
Substituting back into Equation 7.13a produces
$\Bigl[f(t)\varphi(t)\Bigr]_{-\infty}^{\infty}=\int_{-\infty}^{\infty}f^{\prime}(t)\varphi(t)\,dt+\int_{-\infty}^{\infty}f(t)\varphi^{\prime}(t)\,dt$ (7.13b)
Evaluate the boundaries at -$\infty$ and $\infty$. Since $\varphi(t)\in C_{c}^{\infty}(\mathbb{R})$, there has to exist some number $N_{+}$ such that $\varphi(t)=0$ for all $|t|>N_{+}$. Similarly, there must exist some $N_{-}$ such that $\varphi(t)=0$ for all $|t|<N_{-}$. This means that in the limit at $t\rightarrow\pm\infty$ the boundary term is
$\Bigl[f(t)\varphi(t)\Bigr]_{-\infty}^{\infty}\lim_{t\to\infty}f(t)\varphi(t)-\lim_{t\to-\infty}f(t)\varphi(t)=0-0=0$
As a result, Equation 7.13b becomes
$0=\int_{-\infty}^{\infty}f^{\prime}(t)\varphi(t)\,dt+\int_{-\infty}^{\infty}f(t)\varphi^{\prime}(t)\,dt$ (7.13c)
which is equivalent to
$\int_{-\infty}^{\infty}f(t)\varphi^{\prime}(t)\,dt=\int_{-\infty}^{\infty}f^{\prime}(t)\varphi(t)\,dt$ (7.13d)
Notice how ther right hand side of Equation 7.13d is (by definition) $\langle T^{\prime}_{f},\varphi\rangle$ (c.f. Equation 7.11). The left hand side is Equation 7.12: $\langle T_{f},\varphi^{\prime}\rangle$. Therefore,
[…]
The distributional derivative is a unique framework that extends differentiation and its operations to encompass both smooth functions and instantaneous events without contradiction. It provides a unified definition of differentiation that not only applies to smooth functions, but also extends, consistently and rigorously, to sets of events where classical differentiation fails. Integration by parts starts with a theorem about functions and ends with a defining identity that holds for all distributions.
This distributional perspective means that neuronal spikes need not be approximated as signals or rates; they can be analyzed directly as event-structured constraints on time. By formalizing the interaction between event timing and continuous dynamics, this framework provides a mathematically rigorous structure for studying downstream computational properties such as causality, refractoriness, timing sensitivity, and the interplay with geometry in neural systems, concepts that are foundational to understanding neurophysiology and neural dynamics.